### v2

@@ Abstract
Despite advances in learning-based methods, finding valid Lyapunov functions for nonlinear dynamical systems remains challenging. Current neural network approaches face two main issues: challenges in scalable verification and limited interpretability. To address these, we propose an end-to-end framework using transformers to construct analytical Lyapunov functions (local), which simplifies formal verification, enhances interpretability, and provides valuable insights for control engineers. Our framework consists of a transformer-based trainer that generates candidate Lyapunov functions and a falsifier that verifies candidate expressions and refines the model via risk-seeking policy gradient. Unlike Alfarano et al. (2024), which utilizes pre-training and seeks global Lyapunov functions for low-dimensional systems, our model is trained from scratch via reinforcement learning (RL) and succeeds in finding local Lyapunov functions for high-dimensional and non-polynomial systems. Given the analytical nature of the candidates, we employ efficient optimization methods for falsification during training and formal verification tools for the final verification. We demonstrate the efficiency of our approach on a range of nonlinear dynamical systems with up to ten dimensions and show that it can discover Lyapunov functions not previously identified in the control literature.
@@ 1 Introduction
A Lyapunov function is an energy-based certificate function that decreases over time, which is indispensable for stability verification, especially for nonlinear dynamics Khalil (2002). These functions can also offer valuable insights into system behaviors, aiding in controller design. However, designing a Lyapunov function for nonlinear systems has long been considered more of an ‘art’ than a science, even for stable dynamics, due to its inherent complexities. Motivated by this challenge, we have witnessed great interest in the development of computational algorithms for Lyapunov function construction. McGough et al. (2010) employed an evolutionary algorithm for the symbolic computation of Lyapunov functions, but the exponential growth search space impedes its scalability. Alternatively, sum-of-squares (SOS) methods, as discussed in Ahmadi & Majumdar (2016); Dai & Permenter (2023), formulated this task via semi-definite programming (SDP), ensuring stability through candidate polynomial functions. Despite the progress of SOS methods, several theoretical results (Ahmadi et al., 2011; Ahmadi, 2012; Ahmadi & El Khadir, 2018) on asymptotic Lyapunov stability show that even some very simple globally asymptotically stable dynamical systems may not agree with a polynomial Lyapunov function of any degree. In addition, SOS methods suffer numerical sensitivity issues in practice.
@@
Recent advances in deep learning have enabled data-driven neural Lyapunov function with formal verification (Chang et al., 2019; Zhou et al., 2022; Wu et al., 2023; Edwards et al., 2024; Wang et al., 2024; Yang et al., 2024). However, these methods face two key challenges: 1) lack of interpretability and 2) high verification costs Dawson et al. (2023b). Neural networks’ black-box nature limits insights into the system dynamical behavior. Additionally, over-parameterization and nonlinear activations complicate formal verification, leading to scalability issues. Tools like SMT Chang et al. (2019), MIP Wu et al. (2023), and $\alpha,\beta$-CROWN Yang et al. (2024) require small, specialized networks to ensure feasible verification times.
@@
Compared with neural counterparts, analytical Lyapunov functions offer two distinct advantages. First, their symbolic nature offers interpretability, and provides insights for designing stability-guaranteed control policy (Sontag, 1989; Feng et al., 2023a; b; Cui et al., 2023b). Also, they can serve as templates for other systems with similar formulations and potentially incorporate human expert’s feedback in the loop. Second, analytical functions enable efficient verification due to their simplicity and symbolic structure, which seamlessly integrate with formal verification tools like SMT solvers. This reduces parameter complexity and eliminates the need to extract symbolic components or manage complex neural network elements like nonlinear activations. Figure 1 highlights the efficiency of verifying analytical expressions compared to neural networks for a 6-D polynomial system.
@@
Figure 1: We present the statistics of runtime for single SMT verification (precision $e^{-6}$, tolerance error $e^{-1}$) on the 6-D system (Appendix F.4). Each method is validated once per epoch. The red line represents the median. The mean verification time is averaged over one complete training and the value is displayed in bold. Neural Network in FOSSIL 2.0 (Edwards et al., 2024) has a 10-D hidden layer, and Augmented Neural Lyapunov Control (ANLC) (Grande et al., 2023) has two 10D hidden layers. Our formulas have a complexity (number of tokens) of 20 or fewer. For ours, the best candidate in each epoch is verified in this motivation example.
@@
To tackle this challenge, we introduce an end-to-end framework designed to find analytical Lyapunov functions for nonlinear dynamical systems given in analytical form. Building on the transformer’s ability to model complex dependencies Vaswani et al. (2017) and the success of deep symbolic regression methods Holt et al. (2023), our framework deploys a symbolic transformer Kamienny et al. (2022) for Lyapunov function discovery. The transformer’s encoder captures the structural characteristics of system dynamics, while the decoder generates candidate Lyapunov functions by modeling symbolic token distributions. Given the lack of high-quality (local) Lyapunov function datasets, particularly for high-dimensional systems, we propose a reinforcement learning (RL)-based approach to search for Lyapunov function on a per system basis, instead of pre-training like Alfarano et al. (2024). We verify Lyapunov conditions by localized sampling in the neighborhoods of minimizers of the candidate expressions, which are most likely to have violations. The identified counterexamples are then incorporated into the training set for further optimization. Our contributions are summarized as follows:
  • We introduce the first RL framework for directly discovering analytical Lyapunov functions in a per-system manner, bypassing the need for supervised learning with large-scale datasets.
  • We propose a novel and efficient policy optimization pipeline integrating global-optimization-based Lyapunov verification, reward design for candidate Lyapunov evaluation, and risk-seeking policy gradient to optimize the transformer, trainable on machines with limited computation resources.
  • We demonstrate the efficiency of our method on various systems, including non-polynomial dynamics like the pendulum, quadrotor, and power system frequency control. Notably, our approach scales to a 10-D system and discovers a valid local Lyapunov function for power system frequency control with lossy transmission lines, that is previously unknown in the literature.
@@ 2.1 Learning-based Lyapunov Function Construction
In contrast to neural Lyapunov functions, Feng et al. (2024) and Alfarano et al. (2024) derived analytical Lyapunov functions. Feng et al. (2024) combined a neural network with the symbolic regression package PySR (Cranmer, 2023), which approximates the network to produce analytical Lyapunov functions, but the lack of interaction between system dynamics and symbolic regression component limits its potential. Alfarano et al. (2024) pre-trained a transformer on backward- and forward-generated global Lyapunov function datasets, relying on beam search for candidate generation. However, their method cannot adaptively refine the candidate Lyapunov functions if the beam search fails on specific dynamics, and it requires a dataset that is expensive to generate (e.g., thousands of CPU hours for a 5-D dynamics dataset) to achieve adequate generalization during inference. Furthermore, its emphasis on global stability limits its applicability to real-world, nonpolynomial control systems, which typically only admit local stability. Consequently, an RL-based approach that directly searches Lyapunov functions for a given system is indispensable.
@@ 2.2 Symbolic Regression with Generative Model
Symbolic regression is a supervised learning task that seeks to discover an analytical function $f:\mathbb{R}^{n}\to\mathbb{R}$ which approximates the output $y_{i}\in\mathbb{R}$ from input $x_{i}\in\mathbb{R}^{n}$.
RL-based Symbolic Regression. RL-based symbolic regression algorithms (Petersen et al., 2020; Costa et al., 2020; Landajuela et al., 2021) employed generative models, typically RNNs, to generate distributions of symbolic tokens representing mathematical operations and variables, from which analytical expressions are sampled. Rewards, evaluating the quality of the sampled expressions, are measured by fitness metrics like RMSE. Due to the non-differentiable step of converting token sequences into symbolic equations, policy gradients are used to optimize the output distributions. Mundhenk et al. (2021) extended this approach by integrating Genetic Programming (GP) to refine generated expressions, improving framework’s overall performance.
Pre-trained Symbolic Regression methods. These methods are inspired by the success of transformers in Natural Language Processing (NLP) tasks. These algorithms contain two steps: 1) pre-train an encoder-decoder network to model $p(f|\mathcal{D})$ on curated datasets by cross-entropy loss, and 2) sample from this distribution during inference via beam search or Monte Carlo Tree Search (MCTS). Methods like (Biggio et al., 2021; Kamienny et al., 2022; Bendinelli et al., 2023) rely on beam search without gradient refinement, often yielding suboptimal results for out-of-distribution data. In contrast, Holt et al. (2023) integrates RL-based policy gradient optimization with end-to-end RMSE loss for both pre-training and inference, allowing gradient refinement for unseen datasets during inference. Further, (Shojaee et al., 2023; Kamienny et al., 2023) enhance decoding process (expression generation) by incorporating MCTS with feedbacks, such as fitting accuracy and equation complexity.
@@ 3.1 Lyapunov Stability
where $f:\mathcal{D}\to\mathbb{R}^{n}$ is a Lipschitz-continuous vector field, and $\mathcal{D}\subseteq\mathbb{R}^{n}$ is an open set containing the origin that defines the state space. Each $x(t)\in\mathcal{D}$ is a state vector.
@@ 4 Proposed Framework
Figure 2: Framework overview: The transformer takes embeddings of the dynamical system model as input and generates candidate Lyapunov functions in an autoregressive manner. The output is the pre-order traversal of the expression’s binary tree. Candidates are verified using a global-optimization-based verification process, with counterexamples added to the training set for reward calculation. The transformer is updated via the risk-seeking policy gradient. The program terminates once a valid candidate expression is found.
@@ 4.1 Candidate Lyapunov Function Generation from Symbolic Transformer
Expression Representation. Inspired by the deep symbolic regression frameworks, we use a symbolic transformer model as the backbone. The transformer takes a dynamical system $f(x)$ as input and generates candidate analytical Lyapunov functions $\tilde{V}_{\phi}$ such that: $\tilde{V}_{\phi}(x)>0\text{ and }L_{f}\tilde{V}_{\phi}<0,\forall\>x\in\mathcal{D}\backslash\{0\}$. Following the expression representation rules in Lample & Charton (2020), our framework represents symbolic transformer models’ input and output as sequences of symbolic tokens. Each mathematical expression can be converted into a symbolic expression tree, a binary tree where internal nodes are symbolic operators and terminal nodes (leaves in the tree) are variables or constants. Symbolic operators can be either unary (i.e. one child), such as $\sin,\cos$, or binary (i.e. two children), such as +,×+$\times$+ , ×. Furthermore, each symbolic expression tree can be represented as a sequence of node values, either symbolic tokens or numerical coefficients, by its pre-order traversal (i.e. first visiting the parent, then traversing the left child and right child). In this way, each expression obtains a pre-order traversal representation, which can uniquely reconstruct the original expression Petersen et al. (2020). We denote $\tilde{V}_{\phi_{i}}$ as the $i^{\textrm{th}}$ node value in the pre-order traversal of $\tilde{V}_{\phi}$’s expression tree, and $\mathcal{L}_{s}$ as the symbolic library, e.g. {+,×$\{+,\times,\log,\sin,x_{j}\}$, where $\tilde{V}_{\phi_{i}}$ is sampled from.
@@
Dynamics Tokenization. By Definition 3.1, the symbolic transformer models’ input $f(x)$ is composed by $n$ ordinary differential equations $\frac{dx_{i}}{dt}=f_{i}(x)$ for $i=1,\cdots,n$. Each analytical expression $f_{i}(x)$ can be represented as a sequence of symbolic tokens and numerical coefficients by the pre-order traversal of its expression tree. Concatenated the sequences of pre-order traversal for all $f_{i}(x)$, with SOS (start token) and EOS (end token) as separators, we obtain the tokenized dynamics, which is fed into the encoder of symbolic transformer and encoded as a latent vector $\mathcal{F}\in\mathbb{R}^{p}$. The numerical coefficients are tokenized in two schemes: an integer is represented as a sequence of digits in base $b=10$ (e.g. 123 is tokenized as $[1,2,3]$), and a real number is represented in scientific notation rounded to 4 significant digits (e.g. 3.14 is tokenized as $[3,1,4,0,10^{0}]$). A detailed example is illustrated in Figure 3, Appendix A, where we presents the symbolic representations of the simple pendulum dynamics in sequences of pre-order traversal.
@@
Candidate Expression Generation. The decoder generates candidate expressions $\tilde{V}_{\phi}$ in an auto-regressive manner. That is, each token $\tilde{V}_{\phi_{i}}$ in the pre-order traversal of $\tilde{V}_{\phi}$ is sampled from the symbolic library $\mathcal{L}_{s}$ according to conditional distribution $p(\tilde{V}_{\phi_{i}}|\tilde{V}_{\phi_{1:(i-1)}},\phi,f(x))$, where $\tilde{V}_{\phi_{1:(i-1)}}$ is the first $(i-1)$ selected symbolic tokens in the pre-order traversal of $\tilde{V}_{\phi}$. This conditional dependence can be achieved by the decoder which emits a probability distribution $\psi$ over symbolic library $\mathcal{L}_{s}$, conditioned on the previously selected tokens and input dynamics. Since analytical expression in its pre-order traversal is inherently hierarchical, we also deploy the hierarchical tree state representation (Petersen et al., 2020; Holt et al., 2023). This method enhances the decoder input by concatenating the representations of the parent and sibling nodes with previously selected outputs and the dynamics. Once the token sampling process for $\tilde{V}_{\phi}$ is complete, we evaluate the function value at origin $\tilde{V}_{\phi}(0)$ and subtract it from $\tilde{V}_{\phi}$, ensuring the Lyapunov condition $\tilde{V}_{\phi}(0)=0$. Through this process, we can sample a batch of $Q$ candidates $\tilde{\mathcal{V}}_{\phi}=\{\tilde{V}_{\phi}^{i}\sim p(\tilde{V}_{\phi}|\phi,f(x))\}_{i=1}^{Q}$, which will be verified according to the Lyapunov conditions.

### v3

@@ Abstract
Despite advances in learning-based methods, finding valid Lyapunov functions for nonlinear dynamical systems remains challenging. Current neural network approaches face two main issues: challenges in scalable verification and limited interpretability. To address these, we propose an end-to-end framework using transformers to construct analytical Lyapunov functions (local), which simplifies formal verification, enhances interpretability, and provides valuable insights for control engineers. Our framework consists of a transformer-based trainer that generates candidate Lyapunov functions and a falsifier that verifies candidate expressions and refines the model via risk-seeking policy gradient. Unlike Alfarano et al. (2024), which utilizes pre-training and seeks global Lyapunov functions for low-dimensional systems, our model is trained from scratch via reinforcement learning (RL) and succeeds in finding local Lyapunov functions for high-dimensional and non-polynomial systems. Given the symbolic nature of the Lyapunov function candidates, we employ efficient optimization methods for falsification during training and formal verification tools for the final verification. We demonstrate the efficiency of our approach on a range of nonlinear dynamical systems with up to ten dimensions and show that it can discover Lyapunov functions not previously identified in the control literature. Full implementation is available on Github.
@@ 1 Introduction
A Lyapunov function is an energy-like function used to certify stability of dynamical systems. A sufficient condition for stability is that the Lyapunov function decreases along system trajectories. Lyapunov functions also play a central role in controller design, providing formal guarantees of closed-loop stability and robustness Khalil (2002). However, designing a Lyapunov function for nonlinear systems has long been considered more of an ‘art’ than a science, even for stable dynamics, due to its inherent complexities. Motivated by this challenge, we have witnessed great interest in the development of computational algorithms for Lyapunov function construction. McGough et al. (2010) employed an evolutionary algorithm for the symbolic computation of Lyapunov functions, but the exponential growth search space impedes its scalability. Alternatively, sum-of-squares (SOS) methods reformulate the task as a semidefinite program (SDP) that certifies stability with polynomial candidates (Papachristodoulou & Prajna, 2005a; b; Ahmadi & Majumdar, 2016; Dai & Permenter, 2023). However, handling local constraints or non-polynomial dynamics requires auxiliary variables and extra (in)equality constraints (Papachristodoulou & Prajna, 2005a; b), which leads to scalability issues of the SOS methods for real-world problems. Moreover, the theoretical result Ahmadi et al. (2011) on asymptotic Lyapunov stability shows that even some very simple globally asymptotically stable dynamics may not agree with a polynomial Lyapunov function of any degree.
@@
Recent advances in deep learning have enabled data-driven neural Lyapunov function with formal verification (Chang et al., 2019; Zhou et al., 2022; Wu et al., 2023; Dawson et al., 2023b; Edwards et al., 2024; Wang et al., 2024; Yang et al., 2024). However, these methods face two key challenges: 1) lack of interpretability and 2) high verification costs (Dawson et al., 2023b). Neural networks’ black-box nature limits insights into the system’s dynamical behavior. Additionally, over-parameterization and nonlinear activations complicate formal verification, leading to scalability issues. Tools like SMT Chang et al. (2019), MIP Wu et al. (2023), and $\alpha,\beta$-CROWN Yang et al. (2024) require small, specialized networks to ensure feasible verification times.
@@
Compared with neural Lyapunov functions, analytical Lyapunov functions offer two distinct advantages. First, their symbolic nature offers interpretability and provides insights for designing stability-guaranteed control policy (Sontag, 1989; Feng et al., 2023a; b; 2024b; Cui et al., 2023b). Second, analytical functions enable efficient verification due to their simplicity and symbolic structure, which seamlessly integrate with formal verification tools like SMT solvers. This reduces parameter complexity and eliminates the need to verify complex neural network elements like nonlinear activations. Figure 1 highlights the efficiency of verifying analytical expressions compared to neural networks.
@@
Figure 1: We present the statistics of runtime for single SMT verification (precision $e^{-6}$, tolerance error $e^{-1}$) on the 6-D system (Appendix F.4). Each method is validated once per epoch. The red line represents the median. The mean verification time is averaged over one complete training, and the value is displayed in bold. Neural Network in FOSSIL 2.0 Edwards et al. (2024) has a 10-D hidden layer, and Augmented Neural Lyapunov Control (ANLC Grande et al. (2023)) has two 10D hidden layers. Our formulas have a complexity (number of tokens) of 20 or fewer. For ours, the best candidate in each epoch is verified in this motivation example.
@@
To tackle this challenge, we introduce an end-to-end framework designed to find analytical Lyapunov functions for nonlinear dynamical systems given in analytical form. Building on the transformer’s ability to model complex dependencies Vaswani et al. (2017) and the success of deep symbolic regression methods Holt et al. (2023), our framework deploys a symbolic transformer Kamienny et al. (2022) for Lyapunov function discovery. The transformer’s encoder captures system dynamics represented as token sequences derived from the ordinary differential equations (ODEs), while the decoder generates candidate Lyapunov functions by modeling symbolic token distributions. Given the lack of high-quality (local) Lyapunov function datasets, particularly for high-dimensional systems, we propose a reinforcement learning (RL)-based approach to search for Lyapunov functions on a per-system basis, instead of pre-training like Alfarano et al. (2024). We verify Lyapunov conditions by localized sampling in the neighborhoods of minimizers of the candidate expressions, which are most likely to have violations. The identified counterexamples are then incorporated into the training set for further optimization. Our main contributions are summarized as follows:
  • We introduce the first RL-based framework for directly discovering analytical Lyapunov functions for nonlinear dynamical systems, bypassing the need for supervised learning with large-scale datasets.
  • We propose a novel and efficient policy optimization pipeline integrating global-optimization-based Lyapunov verification, reward design for candidate Lyapunov evaluation, and risk-seeking policy gradient to optimize the symbolic transformer, trainable on machines with limited computation resources.
  • We demonstrate the efficiency of our method on various systems, including non-polynomial dynamics like the pendulum, quadrotor, and power system frequency control. Notably, our approach scales to a 10-D system and discovers a valid local Lyapunov function for power system frequency control with lossy transmission lines, which is previously unknown in the literature.
@@ 2.1 Learning-based Lyapunov Function Construction
In contrast to neural Lyapunov functions, Feng et al. (2024c) and Alfarano et al. (2024) derived analytical Lyapunov functions. Feng et al. (2024c) combined a neural network with the symbolic regression package PySR (Cranmer, 2023), which approximates the network to produce analytical Lyapunov functions, but the lack of interaction between system dynamics and symbolic regression component limits its potential. Alfarano et al. (2024) pre-trained a transformer on backward- and forward-generated global Lyapunov function datasets, relying on beam search for candidate generation. However, their method cannot adaptively refine the candidate Lyapunov functions if the beam search fails on specific dynamics, and it requires a dataset that is expensive to generate (e.g., thousands of CPU hours for a 5-D dynamics dataset) to achieve adequate generalization during inference. Furthermore, its emphasis on global stability limits its applicability to real-world, nonpolynomial control systems, which typically only admit local stability. Consequently, an RL-based approach that directly searches Lyapunov functions for a given system is indispensable.
@@ 2.2 Symbolic Regression with Generative Model
Symbolic regression is a supervised learning task, searching for an analytical function $f:\mathbb{R}^{n}\to\mathbb{R}$ that fits $y_{i}\in\mathbb{R}$ from input $x_{i}\in\mathbb{R}^{n}$ (Petersen et al., 2020). With suitable extensions, it can scale to multi-input–multi-output mappings.
RL-based Symbolic Regression. RL-based symbolic regression algorithms Petersen et al. (2020); Costa et al. (2020); Landajuela et al. (2021) employed generative models, typically RNNs, to generate distributions of symbolic tokens representing mathematical operations and variables, from which analytical expressions are sampled. Rewards, evaluating the quality of the sampled expressions, are measured by fitness metrics like RMSE. Due to the non-differentiable step of converting token sequences into symbolic equations, policy gradients are used to optimize the output distributions. Mundhenk et al. (2021) extended this approach by integrating Genetic Programming (GP) to refine generated expressions, improving the overall performance.
Pre-trained Symbolic Regression methods. These methods are inspired by the success of transformers in Natural Language Processing (NLP) tasks. These algorithms contain two steps: 1) pre-train an encoder-decoder network to model $p(f|\mathcal{D})$ on curated datasets by cross-entropy loss, and 2) sample from this distribution during inference via beam search or Monte Carlo Tree Search (MCTS). Methods like Biggio et al. (2021); Kamienny et al. (2022); Bendinelli et al. (2023) rely on beam search without gradient refinement, often yielding suboptimal results for out-of-distribution data. In contrast, Holt et al. (2023) integrates RL-based policy gradient optimization with end-to-end RMSE loss for both pre-training and inference, allowing gradient refinement for unseen datasets during inference. Further, Shojaee et al. (2023); Kamienny et al. (2023) enhance the decoding process (expression generation) by incorporating MCTS with feedback, such as fitting accuracy and equation complexity.
@@ 3 Preliminary
where $f:\mathcal{D}\to\mathbb{R}^{n}$ is a Lipschitz-continuous vector field, and $\mathcal{D}\subseteq\mathbb{R}^{n}$ is a set containing the origin that defines the state space. Each $x(t)\in\mathcal{D}$ is a state vector.
@@ 4 Proposed Framework
Figure 2: Framework overview: The transformer takes embeddings of the dynamical system model as input and generates candidate Lyapunov functions in an autoregressive manner. Hierarchical information is deployed to enhance the model input. For example, when generating the last token $x_{2}$, its parent is $+$, and its sibling is $x_{1}$. The output is the pre-order traversal of the expression’s binary tree. Candidates are verified using a global-optimization-based verification process, with counterexamples added to the training set for reward calculation. The transformer is updated via the risk-seeking policy gradient. The program terminates once a valid expression is found.
@@ 4.1 Candidate Lyapunov Function Generation from Symbolic Transformer
Expression Representation. Inspired by the deep symbolic regression frameworks, we use a symbolic transformer model as the backbone. The transformer takes a dynamical system $f(x)$ as input and generates candidate analytical Lyapunov functions $\tilde{V}_{\phi}$ such that: $\tilde{V}_{\phi}(x)>0\text{ and }L_{f}\tilde{V}_{\phi}<0,\forall\>x\in\mathcal{D}\backslash\{0\}$. Following the expression representation rules in Lample & Charton (2020), our framework represents symbolic transformer models’ input and output as sequences of symbolic tokens. Each mathematical expression can be converted into a symbolic expression tree, a binary tree where internal nodes are symbolic operators and terminal nodes (leaves in the tree) are variables or constants. Symbolic operators can be either unary (i.e., one child), such as $\sin,\cos$, or binary (i.e., two children), such as +,×+$\times$+ , ×. Furthermore, each symbolic expression tree can be represented as a sequence of node values, either symbolic tokens or numerical coefficients, by its pre-order traversal (i.e., first visiting the parent, then traversing the left child and right child). In this way, each expression obtains a pre-order traversal representation, which can uniquely reconstruct the original expression Petersen et al. (2020). We denote $\tilde{V}_{\phi_{i}}$ as the $i^{\textrm{th}}$ node value in the pre-order traversal of $\tilde{V}_{\phi}$’s expression tree, and $\mathcal{L}_{s}$ as the symbolic library, e.g. {+,×$\{+,\times,\log,\sin,x_{j}\}$, where $\tilde{V}_{\phi_{i}}$ is sampled from.
@@
Dynamics Tokenization. By Definition 3.1, the symbolic transformer models’ input $f(x)$ is composed by $n$ ordinary differential equations $\frac{dx_{i}}{dt}=f_{i}(x)$ for $i=1,\cdots,n$. Each analytical expression $f_{i}(x)$ can be represented as a sequence of symbolic tokens and numerical coefficients by the pre-order traversal of its expression tree. Concatenated the sequences of pre-order traversal for all $f_{i}(x)$, with SOS (start token) and EOS (end token) as separators, we obtain the tokenized dynamics, which is fed into the encoder of symbolic transformer and encoded as a latent vector $\mathcal{F}\in\mathbb{R}^{p}$. The numerical coefficients are tokenized in two schemes: an integer is represented as a sequence of digits in base $b=10$ (e.g. 123 is tokenized as $[1,2,3]$), and a real number is represented in scientific notation rounded to 4 significant digits (e.g. 3.14 is tokenized as $[3,1,4,0,10^{0}]$). A detailed example is illustrated in Figure 3, Appendix A, where we present the symbolic representations of the simple pendulum dynamics in sequences of pre-order traversal.
@@
Candidate Expression Generation. The decoder generates candidate expressions $\tilde{V}_{\phi}$ in an auto-regressive manner. That is, each token $\tilde{V}_{\phi_{i}}$ in the pre-order traversal of $\tilde{V}_{\phi}$ is sampled from the symbolic library $\mathcal{L}_{s}$ according to conditional distribution $p(\tilde{V}_{\phi_{i}}|\tilde{V}_{\phi_{1:(i-1)}},\phi,f(x))$, where $\tilde{V}_{\phi_{1:(i-1)}}$ is the first $(i-1)$ selected symbolic tokens in the pre-order traversal of $\tilde{V}_{\phi}$. This conditional dependence can be achieved by the decoder, which emits a probability distribution $\psi$ over the symbolic library $\mathcal{L}_{s}$, conditioned on the previously selected tokens and input dynamics. Since analytical expression in its pre-order traversal is inherently hierarchical, we also deploy the hierarchical tree state representation (Petersen et al., 2020; Holt et al., 2023). This method enhances the decoder input by concatenating the representations of the parent and sibling nodes with previously selected outputs and the dynamics. Once the token sampling process for $\tilde{V}_{\phi}$ is complete, we evaluate the function value at origin $\tilde{V}_{\phi}(0)$ and subtract it from $\tilde{V}_{\phi}$, ensuring the Lyapunov condition $\tilde{V}_{\phi}(0)=0$. Through this process, we can sample a batch of $Q$ candidates $\tilde{\mathcal{V}}_{\phi}=\{\tilde{V}_{\phi}^{i}\sim p(\tilde{V}_{\phi}|\phi,f(x))\}_{i=1}^{Q}$, which will be verified according to the Lyapunov conditions.